# SafeNav: Safe Path Navigation using Landmark Based Localization in a GPS-denied Environment

### v1

@@ Abstract
In battlefield environments, adversaries frequently disrupt GPS signals, requiring alternative localization and navigation methods. Traditional vision-based approaches like Simultaneous Localization and Mapping (SLAM) and Visual Odometry (VO) involve complex sensor fusion and high computational demand, whereas range-free methods like DV-HOP face accuracy and stability challenges in sparse, dynamic networks. This paper proposes LanBLoc-BMM, a navigation approach using landmark-based localization (LanBLoc) combined with a battlefield-specific motion model (BMM) and Extended Kalman Filter (EKF). Its performance is benchmarked against three state-of-the-art visual localization algorithms integrated with BMM and Bayesian filters, evaluated on synthetic and real-imitated trajectory datasets using metrics including Average Displacement Error (ADE), Final Displacement Error (FDE), and a newly introduced Average Weighted Risk Score (AWRS). LanBLoc-BMM (with EKF) demonstrates superior performance in ADE, FDE, and AWRS on real-imitated datasets. Additionally, two safe navigation methods, SafeNav-CHull and SafeNav-Centroid, are introduced by integrating LanBLoc-BMM(EKF) with a novel Risk-Aware RRT* (RAw-RRT*) algorithm for obstacle avoidance and risk exposure minimization. Simulation results in battlefield scenarios indicate SafeNav-Centroid excels in accuracy, risk exposure, and trajectory efficiency, while SafeNav-CHull provides superior computational speed.
@@
Landmark-based Non-GPS localization, Visual Localization, GPS-Denied Environment, Safe Navigation
@@ I Introduction
In military operations, Safe navigation demands a precise and reliable positioning system [1], path planning, and maneuver decision support systems under adversarial conditions. Thus, positioning algorithms are central to safe navigation, especially on a battlefield, where precise location, timing, and coordination are critical for moving troops and ground vehicles.
Challenges: GPS provides reliable location data; however, its dependency on satellite signals makes it vulnerable to jamming and spoofing in combat scenarios.
Several methodologies have been developed to address localization and navigation challenges for GPS-denied environments, including Simultaneous Localization and Mapping (SLAM) [4], Visual Odometry (VO) [5], and other non-GPS-based techniques like DV-Hop Localization [9]. Although SLAM and VO excel in GPS‑denied pose estimation, their heavy computational demand for high‑resolution 3‑D processing and reliance on multi‑sensor fusion [15] make them not feasible for rapid-response systems in battlefield settings where soldiers use resource-constrained and battery-powered mobile devices.
@@
To address these challenges, [3] proposed LanBLoc, a localization framework for non-GPS battlefield environments using stereo vision and landmark recognition with passive camera sensors. However, it does not address navigation and motion planning based on landmark localization, which is the focus of this work.
@@
This research proposes solutions to efficiently guide Moving Entities (ME) along a safe path in a non-GPS environment using LanBLoc as a location measurement model, ensuring efficient and safe navigation from source to destination.
To achieve our objective, we first define ground motion planning and extend it to a Battlefield Motion Model (BMM) that captures the non-linear dynamics of battlefield terrains, including obstacles and hazards. We then propose LanBLoc-BMM, a navigation approach combining LanBLoc for location measurement, BMM for motion planning, and EKF [13] to handle unmodeled influences of battlefield terrain. The model assumes sufficient detectable landmarks for trilateration [3] throughout the maneuver from source to destination. Additionally, we propose two safety check algorithms, SafeNav-CHull and SafeNav-Centroid, to ensure safe state transitions and controlled maneuvering within path boundaries. Integrated with LanBLoc-BMM(EKF), they enable efficient navigation. To handle obstacles, we introduce Risk-Aware RRT* (RAw-RRT*), a variant of RRT* with a novel risk-aware cost function.
We evaluate the performance of the proposed algorithms using standard metrics such as Average Displacement Error (ADE), Final Displacement Error (FDE), and Percent Error. Additionally, we introduce the novel Average Weighted Risk Score (AWRS) metric that measures the trajectory safety based on deviations from hazard boundaries. The performance contributions of our research work are:
@@
1) The LanBLoc-BMM (EKF) algorithm is evaluated against three state-of-the-art localization algorithms integrated with EKF and PF. LanBLoc-BMM achieves the highest accuracy with ADE/FDE of $0.4972/0.4675$, lowest AWRS of $0.3492$, and minimal runtime of $3.98~{}ms$ on synthetic datasets. For real-imitated trajectory datasets, LanBLoc-BMM (EKF) attains an ADE/FDE of $0.2393/0.3095$, the lowest AWRS of $0.1528$, and the shortest runtime of $14.94~{}ms$.
2) Comparing SafeNav-Centroid and SafeNav-CHull algorithms integrated with RAw-RRT*, SafeNav-Centroid demonstrates superior accuracy and safety, with an average percent error improvement of 16.12% over SafeNav-CHull, and ADE/FDE of $1.11/0.99~{}m$ that are $1.67$ times (ADE) and $3.30$ times (FDE) better.
Overall, SafeNav-Centroid generates the shortest path with less exposure to hazard, while SafeNav-Chull provides comparable results with higher computational efficiency.
@@ II Related Works
Recent advances in GPS-free positioning, such as SLAM, VO, and deep learning-based methods, have enhanced autonomous navigation in GPS-denied environments. Vision-based localization uses images or video to estimate a robot’s position. SLAM [4] simultaneously builds maps and estimates pose but requires multi-sensor fusion (e.g., cameras, lidar). VO [6] tracks motion by analyzing frame-to-frame changes. Deep learning approaches [7], especially CNNs, have shown high accuracy across diverse settings.
DV-hop algorithms [9], popular in wireless sensor networks for their simplicity and low-cost hardware, estimate node positions based on hop counts to anchors. However, they perform poorly in dynamic or non-uniform networks due to sensitivity to node mobility.
For filtering, EKF [13] handles nonlinear systems through linearization and is well-suited for localization in battlefield environments due to its efficiency and robustness. Particle Filters (PFs) [14], while better for non-Gaussian systems, require careful tuning of particle count and resampling, impacting performance and computational load.
@@ III Safe Path Navigation Framework
Our system is designed to guide ME through a safe path starting from its current position to the destination. In the context of battlefield navigation, a safe path refers to a field that avoids obstacles and possible hazards. A continuous function that maps a parameter $s$ representing a sequence to a point in the navigable space, such that the path avoids hazards, and collisions with obstacles, and respects the constraints of the navigable space as defined in Eqn. 1.
@@
where $s$ is the parameter, representing sequence, with $s_{0}$ and $s_{f}$ being the start and end sequence numbers, $\mathbb{R}^{n}$ represents the n-dimensional navigable space, $\mathbf{p}(s)$ is the position in the navigable space. The path starts at the initial position $p_{0}$ and ends at the target position $p_{f}$, given by Eqn. 2.
@@
The Control Center (CoC) generates the safe path using the ME’s current state and environmental obstacles or hazards. Fig. 1 shows an overview of our proposed navigation framework for GPS-denied battlefield environments. The ME uses BMM & EKF for motion planning, LanBloc & IMU for state measurement, and SafeNav-Chull or SafeNav-Centroid for heading and maneuvering decisions. The ME initially obtains the safe path as a series of path segments $(s_{1},s_{2}\ldots s_{n})$ each containing trajectory points $p_{i}=\{(x_{i1},y_{i1}),(x_{i2},y_{i2}),\ldots,(x_{im},y_{im})\}$ from the CoC via secure communication. Each path segment $s_{i}=(p_{1},p_{2},\ldots,p_{m})$ uses safety check algorithms to compute control commands. The process begins with the initial State $(x_{i},y_{i},v_{i},{\theta}_{i})$ which includes the initial position coordinates $(x_{i},y_{i})$ , speed $(v_{i})$ and heading angle $({\theta}_{i})$ of the moving entity. The ME runs the stereo vision and YOLOv11-based localization algorithm $(LanBLoc)$ to obtain its initial position and assumes to establish communication with the CoC via a secure communication protocol. However, the architecture of node communication with a CoC is not covered in this research work.
@@ IV-A Landmark-based localization (LanBLoc) of moving object
To address the challenges of deploying physical anchor nodes, recent research presents an alternative solution by recognizing the geographical landmarks as reference points using the YOLOv8-based landmark recognition model[3]. We improve this recognition model using YOLOv11s [8] to identify and precisely locate the reference landmarks. We utilized the pre-trained backbone of YOLOv11s and added a custom classifier head for landmark classification and bounding box coordinates. We utilized Generalized Intersection over Union (GIoU) for bounding box loss, Cross-entropy (CE) for landmark classification loss, and Binary cross-entropy (BCE) for the objectness score. The model was trained for $100$ epochs with a batch size of $16$ using $SGD$ optimizer on the real-world landmark dataset (MSTLandmarkv1) [2] containing 34 landmark classes.
The landmark recognition model achieved high performance for training and validation. For training, it demonstrated a Box Loss Precision (Pr) of 0.938, Recall (R) of 0.963, and mAP @ 0.5 IoU of 0.968, with mAP @[0.5:0.95] IoU at 0.841. Validation results were similarly strong, with a Precision of 0.931, a Recall of 0.953, mAP @ 0.5 IoU of 0.964, and mAP @[0.5:0.95] IoU at 0.834 across all landmark classes.
The distances between the moving entity and the identified landmark anchors are calculated using an efficient stereo-matching technique [2] with high accuracy. Using the localized landmark anchors and the calculated distances to the anchors, the LanBLoc algorithm computes the position of a moving object with an average RMSE of 0.0142m in the X-position and 0.039m in the Y-position.
@@ IV-B Motion Planning of Moving Entity in the Battlefield
Motion planning in battlefield scenarios must account for both physical kinematics and tactical constraints. Our Battlefield Motion Model (BMM) captures key dynamics such as velocity, acceleration, deceleration, and maneuverability (max heading rate) to predict the next state from a given control input. Using update equations 3, 4, and 5, BMM models unit motion through position $\vec{p}t=[x_{t},y_{t}]$, velocity $v_{t}$, and heading $\theta_{t}$, adjusted by desired speed $v{\text{desired}}$, heading change $\Delta\theta$, acceleration $a$, deceleration $d$, and maneuverability $m$, updated at each time step $\Delta t$. Now, the motion model can then be expressed as:
IV-B1 Velocity Update

### v2

@@ Abstract
In a battlefield, adversaries often disrupt or deceive GPS signals, necessitating alternative methods for the localization and safe navigation of troops. Vision-based localization approaches like Simultaneous Localization and Mapping (SLAM) and Visual Odometry (VO) use sensor fusion and are computationally intensive. Range-free localization methods such as DV-HOP suffer from accuracy and stability in dynamic and sparse network topology. Using a landmark-based localization (LanBLoc) and battlefield-specific motion model (BMM) aided with an Extended Kalman Filter (EKF), this paper first proposes a navigation approach called LanBLoc-BMM. The efficiency of the proposed approach was compared with three other state-of-the-art (SOTA) visual localization algorithms integrated with BMM and Bayesian filters EKF & PF for navigation on both synthetic and real trajectory datasets. LanBLoc-BMM (with EKF) shows the best accuracy in terms of Average Displacement Error (ADE) and Final Displacement error (FDE) of 0.2393 and 0.3095, respectively, AWRS of 0.1528 and runtime of 14.94ms on real-dataset, among all SOTA algorithms. Then, two different safe path navigation algorithms called SafeNav-CHull and SafeNav-Centroid are proposed by integrating LanBLoc-BMM( EKF). The SafeNav-Centroid relies on the centroid of the safe path segments, while SafeNav-CHull uses the convex hull of the path segments for movement decisions. Both algorithms integrate a novel Risk-Aware RRT*(RAw-RRT*) algorithm for obstacle avoidance and risk exposure minimization. A new evaluation metric called Average Weighted Risk Score (AWRS) is introduced to evaluate the risk exposure of the trajectory. We presented a simulated battlefield scenario for safe path navigation and evaluated the estimated trajectories of proposed safe path navigation algorithms regarding displacement errors, risk exposure, computation time, and trajectory length. The SafeNav-Centroid shows efficient performances in terms of ADE, FDE, AWRS, and Average percent error, while SafeNav-CHull presents the best computation time performance.
@@
Landmark-based Non-GPS localization, GPS-Denied Environment, Safe Navigation
@@ I Introduction
In military operations, safe navigation demands a precise and reliable positioning system[1], path planning, and maneuver decision support systems under adversarial conditions. Thus, positioning algorithms are central to safe navigation, especially on a battlefield, where precise location, timing, and coordination are critical for moving troops and ground vehicles. In scenarios where adversarial threats like GPS jamming, terrain obstacles, and hazards are prevalent, robust navigation systems play a crucial role in mission planning and maintaining operational effectiveness. Moreover, as autonomous systems like unmanned ground vehicles (UGVs) and unmanned aerial vehicles (UAVs) are becoming integral to modern warfare, the demand for accurate and resilient navigation capabilities continues to grow. These systems ensure the safety, timing, and coordination for successful battlefield maneuvers and mission execution.
Challenges: GPS provides reliable location data however, its dependency on satellite signals makes it vulnerable to jamming and spoofing in combat scenarios[2]. Several methodologies have been developed to address localization and navigation challenges for GPS-Denied environments, including, Simultaneous Localization and Mapping (SLAM) [5], Visual Odometry (VO) [7], and other non-GPS-based techniques like DV-Hop Localization [11]. SLAM technique, where GPS is either unreliable or unavailable, involves the creation of a map of an unknown environment while simultaneously keeping track of the individual’s or vehicle’s location within that environment. VO estimates the position and orientation of a moving object by tracking and analyzing the movement of visual features (such as edges or corners) between consecutive frames of video captured by onboard cameras. While SLAM and VO are powerful techniques for pose estimation and navigation, especially in GPS-denied environments, they come with certain limitations, particularly in safe battlefield navigation. These algorithms require significant power to process high-resolution data or complex 3D environments, which is computationally intensive and might be a constraint in smaller, battery-powered, or rapid-response systems. Also, they require multiple sensor fusion techniques[19], which makes it not feasible in resource-constrained settings like battlefields where soldiers use small and battery-powered mobile devices.
@@
To overcome the challenges and limitations given above, [4] introduced LanBLoc, a localization framework designed for non-GPS battlefield environments. It utilizes stereo vision and landmark recognition, relying solely on passive camera sensors and geographical landmarks as reference points for localization. However, their approach does not tackle navigation and motion planning in non-GPS environments using landmark-based localization, which is the main focus of our work here.
Our Contributions:
Figure 1: A battlefield environment cluttered with landmark clusters, obstacles, and hazard components. The safe path buffer starting from centroid cluster 4 and ending at the centroid of cluster 3 indicates a risk-free zone, guides the moving entity safely in avoiding obstacles, and minimizes risk exposure.
@@
This research primarily focuses on developing solutions to efficiently guide moving military troops, ground vehicles, or reconnaissance units (considered as Moving entities (ME)) along a safe path in a non-GPS environment using LanBLoc as a location measurement model. The safe path refers to a navigational boundary in the battlefield that minimizes the risk exposure due to hazardous components, ensuring the safety, efficiency, and successful navigation of a ME from source to the destination, as illustrated in Fig.1.
To achieve our objective, we first define motion planning of moving objects for ground movement and then adapt it for battlefield scenarios called Battlefield Motion Model (BMM) that captures the essential non-linear dynamics of battlefield terrain, such as obstacles and hazards. Then, we propose a navigation approach LanBLoc-BMM using the LanBLoc as a location measurement model and BMM for motion planning, integrating Extended Kalman Filter (EKF) [17] and Particle Filter (PF) [18] separately to handle unmodeled influences of battlefield terrain. The performance of our proposed LanBLoc-BMM approach is evaluated and compared with state-of-the-art (SOTA) visual localization methods from literature such as DeepLoc-GPS, DeepLoc-EKF-GPS [8], Multiple Sensor Fusion [19], using synthetic and real trajectory datasets. Experimental results showed that the LanBLoc-BMM approach, when integrated with EKF, produces the best navigation performance for both synthetic and real trajectory datasets, demonstrating the effective application of the LanBLoc framework as a location measurement model in battlefield navigation applications. Thus, our LanBLoc-BMM(EKF) based navigation approach helps to navigate the moving entity in a GPS-denied battlefield environment from the source to the target destination, as shown in Fig 1. The proposed navigation approach assumes that the battlefield environment offers sufficient landmarks within the detectable range of moving objects to facilitate trilateration, like in [4]. In the figure, circular shapes indicate the landmark clusters, and the bounded region(strip) starting from the centroid of cluster 4 and ending at cluster 3 indicates the safe path.
Additionally, to ensure the safety of the moving entity, we then proposed two separate algorithms called SafeNav-CHull and SafeNav-Centroid that ensure safety checks of the next possible state and controlled maneuvering within the safe path boundary. Each algorithm is then integrated with a LanBLoc-BMM(EKF) for safe and efficient navigation along the given safe path. To handle the obstacles within the safe path, we also proposed a new variant of the Rapidly Exploring Random Tree-Star (RRT*) algorithm called Risk-Aware RRT* (RAw-RRT*) by defining a novel risk-aware cost function. The SafeNav-Centroid computes the centroids of the given safe path segments for controlled maneuvering of the units, ensuring safety and the shortest trajectory. On the other hand, the SafeNav-CHull applies a point-in-polygon test [15] on the convex hull [13] of safe path segments to ensure the safety and control input calculation. RAw-RRT* ensures an obstacle-free path with the shortest length and minimal risk of hazards. We evaluated the performance of the proposed algorithms using standard metrics such as Average Displacement Errors(ADE), Final Displacement Error, and Percent Error. Additionally, to evaluate the safety, we defined a metric called Average weighted Risk Score (AWRS), which measures the average risk exposure of trajectory generated by navigation algorithms by taking account of the deviation of trajectory points from the hazard boundary. The performance contributions of our research work are:
@@
1) We evaluate and compare the navigation efficiency of the LanBLoc-BMM algorithm with three other SOTA localization algorithms integrated with EKF and PF for navigation on both synthetic and real trajectory datasets. We observe the best accuracy $ADE/FDE$ of 0.4972 / 0.4675, minimal AWRS of 0.3492, and minimal runtime of 3.98ms on synthetic datasets and achieved ($ADE/FDE$ of 0.2393 / 0.3095), minimal AWRS of 0.1528 and the least runtime of 14.94 ms for LanBLoc-BMM (EKF) on a real dataset.
2) We compare the performance of proposed SafeNav-CHull and SafeNav-Centroid algorithms, both integrated with RAw-RRT*. We obtain an average percent error in trajectory estimation of 5.46% for the SafeNav-Centroid algorithm which is a 16.12% improvement over the SafeNav-CHull approach. We obtain an average ADE/FDE of 1.11/0.99m, for the SafeNav-Centroid algorithm, which is 1.67 times and 3.30 times better, respectively, than the ADE and FDE of the SafeNav-CHull algorithm. SafeNav-Centroid also presents lower AWRS compared to SafeNav-Chull; however, SafeNav-Chull demonstrates higher efficiency in terms of computational time. Overall, SaveNav-Centroid generates the shortest path with less exposure to hazard, while SafeNav-Chull provides comparable results with higher computational efficiency.
@@ II Related Works
The recent advancements in GPS-free positioning, including SLAM and VO, have significantly improved the navigation and mapping capabilities of autonomous systems and unmanned ground vehicles. Vision-based methods[5] provide precise and reliable localization results in GPS-denied environments. They utilize image and video data to determine the robot’s location within its surroundings. SLAM technique  [5] estimates the position and orientation $(pose)$ of a robot within an environment while simultaneously creating its map. However, this technique requires a fusion of multiple cameras, lidar, or radar sensors to generate a map of its surroundings. VO methods [8] also estimate the robot’s motion by tracking the changes in images or video frames captured by the robot. Deep Learning-based Localization[9], particularly convolutional neural networks (CNNs), have demonstrated impressive performance in vision-based localization in various environments. DV-hop-based range-free localization algorithms commonly used in wireless sensor networks $(WSNs)$[12] have received enormous attention due to their low-cost hardware requirements for nodes and simplicity in implementation. The algorithm estimates the location of unknown nodes based on their hop count to known anchor nodes and the known location of the anchor nodes. However, these methods only perform well in static, isotropic, and uniformly distributed network topology. They are sensitive to node mobility and cannot adapt to dynamic network conditions. EKF [17] extends the Kalman filter to handle nonlinear systems by linearizing them at each time step by calculating Jacobians of estimating functions. PFs though suitable for non-Gaussian and highly nonlinear systems[18] can be computationally expensive if the right number of particles is not selected. So, careful tuning of the right number of particles and the resampling strategy is required, which can eventually affect the filter’s performance. The $EKF$ is effective for battlefield scenarios due to its balance of accuracy, computational efficiency, robustness, and ability to handle non-linear systems, making it widely used in military applications.
@@ III Safe Path Navigation Framework
Our system is designed to guide a mobile troop through an obstacle and hazard-free trajectory starting from its current position to the destination. The moving entity (node) is required to navigate within the safe path to ensure the safety, survivability, and success of the mission.
In the context of battlefield navigation, a safe path refers to a field that avoids obstacles and possible hazards. A continuous function that maps a parameter $s$ representing a sequence to a point in the navigable space, such that the path avoids hazards, and collisions with obstacles and respects the constraints of the navigable space. Mathematically, a safe path $\mathbf{p}(s)$ can be defined as follows:
@@
where $s$ is the parameter, representing sequence, with $s_{0}$ and $s_{f}$ being the start and end sequence numbers, $\mathbb{R}^{n}$ represents the n-dimensional navigable space, $\mathbf{p}(s)$ is the position in the navigable space. The path starts at the initial position $p_{0}$ and ends at the target position $pf$, which is given by:
@@
The safe path is generated by the control center based on the current state of the moving node and obstacle/hazard components cluttered in the region of interest. The node initially obtains the safe path as a series of path segments $(s_{1},s_{2}\ldots s_{n})$ each containing trajectory points $pi=\{(x_{i1},y_{i1}),(x_{i2},y_{i2}),\ldots,(x_{im},y_{im})\}$ from the command and control center via secure communication. For each path segment $s_{i}=(p_{1},p_{2},\ldots p_{m})$, the node can use either SafeNav-Chull or SafeNav-Controid algorithms to compute control commands. Fig. 2 shows an overview of our proposed navigation framework for GPS-denied battlefield environments. The moving entity uses a BMM & EKF for motion planning, LanBloc & IMU for state measurement, and SafeNav-Chull or SafeNav-Controid for heading and maneuvering decisions. The process begins with the initial State $(x_{i},y_{i},v_{i},{\theta}_{i})$ which includes the initial position coordinates $(x_{i},y_{i})$ , speed $(v_{i})$ and heading angle $({\theta}_{i})$ of the moving entity. The moving entity runs the stereo vision and YOLOv11-based localization algorithm $(LanBLoc)$ to obtain its initial position and assumes to establish communication with the control center via a secure communication protocol. However, the architecture of node communication with a control center is not covered in this research work.
@@ IV-A Landmark-based localization (LanBLoc) of moving object
To address the challenges of deploying physical anchor nodes, recent research presents a solution by recognizing the geographical landmarks as reference points using the YOLOv8-based landmark recognition model[4]. We improve this landmark recognition model using YOLOv11s [10] to identify and precisely locate the reference landmarks. We utilized the pre-trained backbone of YOLOv11s and added a custom classifier head for landmark classification and bounding box coordinates. We utilized Generalized Intersection over Union (GIoU) for bounding box loss, Cross-entropy (CE) for landmark classification loss, and Binary cross-entropy (BCE) for the objectness score. The model was trained for $100$ epochs with a batch size of $16$ using $SGD$ optimizer on the real-world landmark dataset (MSTLandmarkv1) [3] containing 34 landmark classes. We applied pre-processing and augmentation on the image data before training by flipping, rotating, scaling, noising, and lighting variations to make the landmark dataset more realistic to battlefield scenarios and robust to variations in input data.
The landmark recognition model achieved high performance for training and validation. For training, it demonstrated a Box Loss Precision (Pr) of 0.938, Recall (R) of 0.963, and mAP @ 0.5 IoU of 0.968, with mAP @[0.5:0.95] IoU at 0.841. Validation results were similarly strong, with a Precision of 0.931, Recall of 0.953, mAP @ 0.5 IoU of 0.964, and mAP @[0.5:0.95] IoU at 0.834 across all landmark classes.
The distances between the moving entity and the identified landmark anchors are calculated using an efficient stereo-matching technique [3] with high accuracy. Using the localized landmark anchors and the calculated distances to the anchors, the LanBLoc algorithm computes the position of a moving object with an average RMSE of 0.0142m in the X-position and 0.039m in the Y-position. By utilizing the recognized landmarks as anchor nodes for trilateration, we enhanced the accuracy and reliability of node localization with further position optimization.
@@ IV-B Motion Planning of Moving Entity in the Battlefield
Developing a motion planning of a moving entity for a battlefield scenario involves considering both the physical aspects of motion and the tactical constraints imposed by the battlefield. Our Battlefield Motion Model (BMM) incorporates kinematic aspects of the vehicle or troop unit such as velocity, acceleration, and deceleration limits, and its maneuverability, which indicates the maximum rate of change of heading. Given a control command and current state, a predicting function updates the unit’s next state based on those factors. The BMM is represented in Eqns: 3,4, and 5. The parameters are given in the table I.
TABLE I: Parameters used in the Battlefield Motion Model
Parameters Description
$\vec{p}_{t}=[x_{t},y_{t}]$ Position vector at time $t$
$v_{t}$ Velocity at time $t$
$\theta_{t}$ Heading (direction) at time $t$
$v_{\text{desired}}$ Desired velocity (speed command)
$\Delta\theta$ Change in heading (heading command)
$a$ Acceleration capability of the unit
$d$ Deceleration capability of the unit
$m$ Maneuverability (maximum rate of change of heading)
$\Delta t$ Time step
Now, the motion model can then be expressed as:
IV-B1 Velocity Update: